OpenStudy (anonymous):

The set S consists of all odd multiples of 3: S = {..., -9, -3, 3, 9}. If the integers x and y are in S, then which of the following must also be in S? a.) xy b.) x +y c.) x-y d.) x/y e.) -x - y

OpenStudy (fellowroot):

From advanced cal. If S is odd multiples of 3 then... S={ 3(2n+1)}, where n is an integer also x and y belong to S let y= 2n+1, x=2m+1 for a.) xy 3(2n+1)(3)(2m+1) 9(2n+1)(2m+1) 9(4nm+2n+2m+1) 9(2)[(nm+n+m)+1] 9(2k+1), where k =nm+n+m, it also does not matter if k is even or odd since (2k+1) will always be odd so... 9(2k+1) 3(3)(2k+1) 3(6k+3) end of proof... 6k+3 is always odd and it is a multiple of 3, it meets the requirments now for b.) x+y 3(2n+1) + 3(2m+1) 3[(2n+1)+(2m+1)] 3[2m+2n+2] 3[(2)(m+n+1) where m+n=k 3[2k+1] 2k+1 is always odd and its being multiplied by 3 end of proof. use the same idea with the c, d, ect.